Five friends have been borrowing and lending money to each other. John owes Mary $1, Tom owes Ken $1, Ken owes Mary $2, Mary owes Lynn $3, Tom owes Lynn $5, Lynn owes Ken $1, Tom owes Mary $2, Ken owes Tom $1, and Tom owes
John $3. If they all receive a $10 allowance and pay off all their debts, who will have the most money?
(a) John (b) Mary (c) Tom (d) Ken (e) Lynn

Answer by richwmiller(17219)

five friends
m,k,j,t,l
Add money that they owe to the receiver; Subtract money they owe from themselves
j=10 -1+3,
m=10+1+2-3+2,
k=10+1-2+1-1,
l=10+3+5-1,
t=10-1-5-2+1-3
t = 0, l = 17, k = 9, m = 12, j = 12
Who has the most?